OpenStudy (anonymous):

Find the axis of symmetry of the parabola formed by the following quadratic function: y = 3x2 + 7x – 4 x = 7/3 x = -7/3 x = 7/6 x = -7/6

OpenStudy (jackellyn):

I think this equation will help you find the axis x = -b/2a

OpenStudy (anonymous):

how do you do it

OpenStudy (jackellyn):

b=7 a=3 plug into the equation x=-7/3(2) x=-7/6